OpenStudy (anonymous):

A right triangle has base (x - 6) units and height (6x - 12) units. Part A: What is the square of the length of the hypotenuse of the triangle? Show your work. (4 points) Part B: What is the area of the triangle? (3 points) Part C: Using the solution obtained in Part B, explain the closure property of multiplication of polynomials. (3 points) I know A but I need help with B

OpenStudy (anonymous):

@Hero @amistre64

OpenStudy (anonymous):

Do you know what the distance formula is?

OpenStudy (anonymous):

I don't need distance, I need the area and Pythagorean Theorem. @agentc0re

OpenStudy (anonymous):

First question asks the square of the length of the hypotenuse. Also do you know the area formula for triangles? 3rdly, the distance formula is pythagoreans forumla

OpenStudy (anonymous):

x^2+y^2=c^2, I already got that part, c^2=37x^2-180=c^2 I need part B and maybe part C. The are for a triangle is 1/2b*h

OpenStudy (anonymous):

@agentc0re

OpenStudy (anonymous):

sorry the site has been laggy

OpenStudy (anonymous):

well b = base and h = height. in this case we know what our base and height are, they are equations. plug those directly in and you'll get an answer for area.

OpenStudy (anonymous):

As for C, i don't know what the closure property is. I'd just google that and find out what it is.